# A Systematic Evaluation of Euclidean Alignment with Deep Learning for EEG Decoding

### v3

@@ 1 Introduction
This is the problem targeted by the field of Transfer Learning (TL) [35], which aims to improve the transferability of models between domains, where each domain can be an individual in a given task. There are two main approaches for TL: task and domain adaptation [29]. In task adaptation, one trains the model in a given domain and, when applying it to the target domains, fine-tune the model parameters to improve the performance on the target domain. In domain adaptation, we use the same model in both domains but change the input data so that both domains have similar representations.
When considered in the BCI application, a task adaptation would be to train the model on another subject and fine-tune the model using labeled data from the target [45, 15]. When using domain adaptation, one would apply pre-processing transformations to make the source and target EEG signals more similar [36, 48]. One well-known technique is the Riemannian Alignments (RA)  [30, 47, 23, 2], which aims to reduce the distance between the subjects’ covariance matrices in the Riemannian manifold [44]. For instance, Rodrigues et al. [30] propose a method based on the Procrustes analysis, applying geometrical transformations (re-centering, re-scale, rotation) on the Riemannian manifold. On the other hand, Zanini et al. [47] insert paradigm information trying to re-centering based on the resting state trials. However, RA transformations are done on covariance matrices, which limits the use of standard state-of-the-art DL models that work on raw EEG data. Even though there are efforts to apply SPDNets [14] to EEG decoding [39], these networks have not yet been well studied for transfer learning tasks or with alignments.
Another domain adaptation method is the Euclidean Alignment (EA) [13], which works directly on the EEG signal. This form of data alignment reduces the distance between data distributions by centering their Euclidean mean covariance matrix to the identity. As a result of its simplicity and computational efficiency, there has been an increased interest in using EA for EEG decoding, usually as a pre-processing step combined with state-of-the-art models [37, 19, 7], and has been showing some good results on both traditional machine learning [9, 22] and deep learning pipelines [41] and in the context of merging datasets  [27]. Despite its growing popularity, there are limited studies on how EA can affect deep learning models’ training and transfer performance for shared models trained with data from multiple subjects and individual models trained with data from single subjects.
@@ 2.9 Model Training
We trained the shared models using the leave-one-subject-out paradigm. We used all data from the source subjects since the models were evaluated in different target subjects. Inside the source subjects set, we initially randomly separated 20% for validation and 80% for training. For the shared models, we used a batch of $64$ trials with $1000$ epochs, an early stopping of $250$ iterations, and a dropout rate of $25\%$. For each architecture, we executed a Grid search using leave-one-group-out cross-validation to optimize the learning rate and weight decay, with the search space shown in Table 1. We ended up with $w_{d}=1e^{-5}$ weight decay for all models and a learning rate of $\eta=1e^{-3}$ for both EEGNet and DeepNet and $\eta=8.25e^{-4}$ for Shallow ConvNet. When using fine-tuning, we retrained them for an additional $600$ epochs, with a patience of $150$. During fine-tuning, we used the first $24$ trials from each training subject and the first $24$ trials from the target.
@@ 3.1 Transfer-learning using Euclidean aligned shared models
Euclidean Alignment improved the average decoding performance on all evaluated datasets and architectures, as shown in line No-EA, offline-EA, and online-EA in Table 2. The offline-EA had an accuracy of 1.26% better than the online-EA, which occurs because all the target data is used to estimate the covariance matrices, resulting in a better alignment. However, the online-EA had an overall mean accuracy of 4.33% better than the non-aligned case, showing the benefit of using EA in a more realistic scenario. Finally, using EA and RA resulted in very similar results, with slightly better accuracies for one of the approaches depending on the model or dataset.
@@ 3.3 Analyzing transferability between subjects
The analysis also indicates that, in general, individuals who are bad donors (sources), i.e., whose models, on average, do not have a good performance on different target subjects, also tend to be bad receivers (targets), and vice versa, as shown in Figure 8. Applying Euclidean Alignment did not change the structure of good/bad donors and receivers but increased the correlation, as can be observed in Table 3. This behavior was expected since, in the aligned case, the data distribution becomes more similar between subjects. If the data from a source subject is useful for the target ones, we should expect the reverse also to occur. Without alignment, data from different subjects are most likely spread in different areas in the domain space.
@@ 3.4 Majority-voting classifiers for transfer learning
It is possible to notice from Table 4 that, when training individual models, using EA also resulted in clear benefits, with a 3.71% increase in the overall mean accuracy when comparing the 3-model classifiers with and without EA. For all datasets and neural network models, the best mean accuracies occurred for the 3 and 5-model classifiers with EA.
@@ 4 Discussion
Fine-tuning resulted in no improvements in accuracy for the shared model with EA and a 1.43% increase in accuracy without EA. The small increase without EA possibly occurred because there was only a very small amount of data to fine-tune the model. Using EA, since the models were aligned from the beginning, this small fine-tuning was not sufficient to improve the accuracy.
We also evaluated the scenario with multiple models trained with data from different subjects. One effect noted in other works [31] applying RA is that subjects with good self-scores tend to be good “receivers". Using EA, we did not evaluate the self-scores, but we noted that subjects are good “donors" tend to be good “receivers" and bad “donors" tend to be bad “receivers." We did not try to train the shared model using only the “good” donors, but this could be a way to improve the transfer accuracies.
When training individual models, using EA also resulted in clear benefits when selecting only the best source model and majority-voting classifiers of 3 and 5 models. Using weighted majority voting from 3 or 5 models worked better than just using the best model but had a 3.6% reduction in the average accuracy compared to the online shared model. One advantage of the majority-voting models is that adding more source subjects to the collection of source models is easy. We only have to add the new source models to the selection process for the best transferable models for each target subject. One disadvantage is that the inference takes a little longer. However, it should be fast enough for real applications.
Kostas and Rudzicz [19] discusses the concepts of domain generalization (DG), where we adapt the source domains (subjects) to be useful for multiple target domains (subjects), such as when using EA, and domain adaptation (DA), where the domains are adapted to a single target subject. An example of DA would be selecting different source subjects to transfer to each target subject, as done in the majority-voting models. The authors argue that one can apply DG and DA concurrently and can improve transfer learning, as we also verified in our experiments with our majority-voting classifiers.
@@ 4.1 Limitations
It is important to address some limitations of our study. Firstly, we didn’t explore the impact of different ranges of hyperparameters, such as learning rate, in models with and without EA. For instance, a model with aligned input may accept larger learning rates since there would be less difference between batches. We also used the same hyperparameters for all individual models, but it is well-known that subject-specific hyperparameters can improve accuracy. Other interesting contributions would be to explore the faster convergence of aligned models better and analyze for improvements in the variance of predictions, using, for example, the Monte-Carlo Dropout technique.
One of the approaches employed in this study was to use a weighted majority voting mechanism over the best individual models. However, combining models optimized for each individual is not the best way to generalize knowledge across diverse subjects. We selected this approach because the focus of the evaluation of individual models was to determine if Euclidean Alignment improves transferability between subjects.
@@ Acknowledgments
BJ and RYC are grateful to São Paulo Research Foundation (FAPESP) for the financial support (grants 22/08920-0 and 23/06407-7). The work of BA was supported in part by the CAPES under Grant 001/003 and Data IA mobilité international, Convergence Institute as part of the “Programme d’Investissement d’Avenir” (ANR-17-CONV-0003) operated by LISN.

### v4

@@ 1 Introduction
This is the problem targeted by the field of Transfer Learning (TL) [40, 2, 15], which aims to improve the transferability of models between domains, where each domain can be an individual in a given task. There are two main approaches for TL: task and domain adaptation [33]. In task adaptation, one trains the model in a given domain and, when applying it to the target domains, fine-tune the model parameters to improve the performance on the target domain. In domain adaptation, we use the same model in both domains but change the input data so that both domains have similar representations.
When considered in the BCI application, a task adaptation would be to train the model on another subject and fine-tune the model using labeled data from the target [50, 18]. When using domain adaptation, one would apply pre-processing transformations to make the source and target EEG signals more similar [41, 53]. One well-known technique is the Riemannian Alignments (RA)  [35, 52, 26, 3], which aims to reduce the distance between the subjects’ covariance matrices in the Riemannian manifold [49]. For instance, Rodrigues et al. [35] propose a method based on the Procrustes analysis, applying geometrical transformations (re-centering, re-scale, rotation) on the Riemannian manifold. On the other hand, Zanini et al. [52] insert paradigm information trying to re-centering based on the resting state trials. However, RA transformations are done on covariance matrices, which are incompatible with standard state-of-the-art DL models that operate on raw EEG data. There have been efforts to apply DL networks that work on covariance matrices, such as SPDNets [17] to EEG decoding [44, 8] and Attention models to detect cognitive impairements [34], but the number of studies is still limited.
Another domain adaptation method is the Euclidean Alignment (EA) [16], which works directly on the EEG signal. This form of data alignment reduces the distance between data distributions by centering their Euclidean mean covariance matrix to the identity. As a result of its simplicity and computational efficiency, there has been an increased interest in using EA for EEG decoding, usually as a pre-processing step combined with state-of-the-art models [42, 22, 9], and has been showing some good results on both traditional machine learning [11, 25] and deep learning pipelines [46] and in the context of merging datasets  [31, 28]. Despite its growing popularity, there are limited studies on how EA can affect deep learning models’ training and transfer performance for shared models trained with data from multiple subjects and individual models trained with data from single subjects.
@@ 2.9 Model Training
We trained the shared models using the leave-one-subject-out paradigm. We used all data from the source subjects since the models were evaluated in different target subjects. Inside the source subjects set, we initially randomly separated 20% for validation and 80% for training. For the shared models, we used a batch of $64$ trials with $1000$ epochs, an early stopping of $250$ iterations, and a dropout rate of $25\%$. For each architecture, we executed a Grid search using leave-one-group-out cross-validation to optimize the learning rate and weight decay, with the search space shown in Table 1. We included a lower learning rate for shared models because they have more diversified data, which could lead to more instability during training. We ended up with $w_{d}=1e^{-5}$ weight decay for all models and a learning rate of $\eta=1e^{-3}$ for both EEGNet and DeepNet and $\eta=8.25e^{-4}$ for Shallow ConvNet. When using fine-tuning, we retrained them for an additional $600$ epochs, with a patience of $150$. During fine-tuning, we used the first $24$ trials from each training subject and the first $24$ trials from the target.
@@ 3.1 Transfer-learning using Euclidean aligned shared models
Euclidean Alignment improved the average decoding performance on all evaluated datasets and architectures, as shown in line No-EA, offline-EA, and online-EA in Table 2. The offline-EA had an accuracy of 1.26% better than the online-EA, which occurs because all the target data is used to estimate the covariance matrices, resulting in a better alignment. However, the online-EA had an overall mean accuracy of 4.33% better than the non-aligned case, showing the benefit of using EA in a more realistic scenario. Finally, using EA and RA resulted in very similar results, with slightly better accuracies for one of the approaches depending on the model or dataset. We use the Euclidean Alignment in the next experiments since it is the most used alignment method when combined with DL models.
@@ 3.3 Analyzing transferability between subjects
The analysis also indicates that, in general, individuals who are bad donors (sources), i.e., whose models, on average, do not have a good performance on different target subjects, also tend to be bad receivers (targets), and vice versa, as shown in Figure 8. Applying Euclidean Alignment did not change the structure of good/bad donors and receivers but increased the correlation, as can be observed in Table 3. This behavior was expected since, in the aligned case, the data distribution becomes more similar between subjects. If the data from a source subject is useful for the target ones, we should expect the reverse to occur as well. Without alignment, data from different subjects are most likely spread in different areas in the domain space.
@@ 3.4 Majority-voting classifiers for transfer learning
It is possible to notice from Table 4 that, when training individual models, using EA also resulted in clear benefits, with a 3.71% increase in the overall mean accuracy when comparing the 3-model classifiers with and without EA. The best mean accuracies occurred for all datasets and neural network models for the 3 and 5-model classifiers with EA.
@@ 4 Discussion
Fine-tuning resulted in no improvements in accuracy for the shared model with EA and a 1.43% increase in accuracy without EA. The small increase without EA possibly occurred because there was only a small amount of data to fine-tune the model. Using EA, since the models were aligned from the beginning, this small fine-tuning was insufficient to improve the accuracy.
We also evaluated the scenario with multiple models trained with data from different subjects. One effect noted in other works applying RA is that subjects with good self-scores tend to be good “receivers" [36]. Using EA, we did not evaluate the self-scores, but we noted that subjects are good “donors" tend to be good “receivers" and bad “donors" tend to be bad “receivers." We did not try to train the shared model using only the “good” donors, but this could be a way to improve the transfer accuracies.
When training individual models, using EA also resulted in clear benefits when selecting only the best source model and majority-voting classifiers of 3 and 5 models. Using weighted majority voting from 3 or 5 models worked better than just using the best model but had a 3.6% reduction in the average accuracy compared to the online shared model. One advantage of the majority-voting models is that adding more source subjects to the collection of source models is easy. We only need to add the new source models to the selection process for the best transferable models for each target subject. One disadvantage is that the inference takes longer. However, it should be fast enough for real applications.
Kostas and Rudzicz [22] discusses the concepts of domain generalization (DG), where we adapt the source domains (subjects) to be useful for multiple target domains (subjects), such as when using EA, and domain adaptation (DA), where the domains are adapted to a single target subject. An example of DA would be selecting different source subjects to transfer to each target subject, as done in the majority-voting models. The authors argue that one can apply DG and DA concurrently and improve transfer learning, as we verified in our experiments with our majority-voting classifiers.
@@ 4.1 Limitations
It is important to address some of the limitations of our study. Firstly, we did not explore the impact of different ranges of hyperparameters, such as learning rate, in models with and without EA. For instance, a model with aligned input may accept larger learning rates since there would be less difference between batches. We also used the same hyperparameters for all individual models, but it is well-known that subject-specific hyperparameters can improve accuracy. Other interesting contributions would be to explore the faster convergence of aligned models better and analyze for improvements in the variance of predictions, using, for example, the Monte-Carlo Dropout technique.
One of the approaches employed in this study was to use a weighted majority voting mechanism over the best individual models. However, there are better ways to generalize knowledge across diverse subjects than combining models optimized for each individual. We selected this approach because the focus of the evaluation of individual models was to determine if Euclidean Alignment improves transferability between subjects.
@@ Acknowledgments
BJ and RYC thank the São Paulo Research Foundation (FAPESP) for the financial support (grants 22/08920-0 and 23/06407-7). The work of BA was supported in part by the CAPES under Grant 001/003 and Data IA mobilité international, Convergence Institute as part of the “Programme d’Investissement d’Avenir” (ANR-17-CONV-0003) operated by LISN.